OpenStudy (anonymous):

5 mathematical examples of the additive inverse, performing any calculations to explain why it is an example of the additive inverse accurately and correctly.

OpenStudy (anonymous):

!@Mehek14 @Freez

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@DARTHVADER2900

OpenStudy (anonymous):

@hhelpplzzzz

OpenStudy (anonymous):

@Thesmarterone can you helpme with this nobody wants to help me :(

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Additive_inverse I am not so good in math, so here is this :)

OpenStudy (phi):

every number has an "additive inverse" it is the number you add to the original so they add up to 0 for example: given 1, its additive inverse is -1, because 1+ -1 = 0 I am sure you can pick a number, and then write its additive inverse (by putting a minus sign in front of it)

OpenStudy (phi):

so pick a number. what number did you pick ?

OpenStudy (anonymous):

5

OpenStudy (phi):

now write down that same number with a minus sign in front.

OpenStudy (anonymous):

5+ -5=0 ??

OpenStudy (phi):

yes, that shows that -5 is an additive inverse of 5 (and also 5 is the additive inverse of -5) now pick another number.

OpenStudy (anonymous):

9+ -9=0

OpenStudy (phi):

yes. not too hard, right? Here is one: -2 what is the additive inverse of -2 ?

OpenStudy (anonymous):

0

OpenStudy (phi):

put a - sign in front of -2, to get -(-2) two minus make a plus, so -(-2) is 2 2 is the additive inverse of -2 and -2 + 2 = 0

OpenStudy (anonymous):

okay so this is what I'm going to do 2+ - 2 = 0 5+ - =0 7+ - = 0 9+ - 9 = 0 12 + -12=0 em I right??

OpenStudy (phi):

the 2nd and 3rd lines left out a number

OpenStudy (anonymous):

oh sorry 7 + -7 = 0 5 + -5 = 0 hey can you helpme withthis Include at least 5 mathematical examples of the multiplicative inverse, performing any calculations to explain why it is an example of the multiplicative inverse accurately and correctly.

OpenStudy (phi):

multiplying has almost the same idea except we 1) multiply (instead of add) 2) when we multiply the two numbers you get 1 (not zero) examples: \[ 2 \text{ and } \frac{1}{2} \\ \frac{2}{3} \text{ and } \frac{3}{2}\] given a number, you find the multiplicative inverse by "flipping" the number

OpenStudy (anonymous):

oh like this 2/3 3/2 4/5 5/4 6/8 8/6 10/12/10/12 like that????

OpenStudy (phi):

and multiply each pair. what do you get ?

OpenStudy (anonymous):

the first on eis 1/1 right????

OpenStudy (phi):

1/1 simplifies to 1

OpenStudy (anonymous):

so for all of them is 1?????

OpenStudy (anonymous):

can oyu help me for the last one Explain the additive and multiplicative inverse properties of numbers accurately.

OpenStudy (phi):

yes. that is how you prove they are multiplicative inverses

OpenStudy (anonymous):

oh okay that was very easy!

OpenStudy (phi):

be sure you can find the multiplicative inverse of 1/4 (for example). any idea ?

OpenStudy (anonymous):

is it one??

OpenStudy (phi):

no. remember you want 1/4 * something = 1 1/4 * 1 is 1/4 to find the multiplicative inverse, "flip" 1/4

OpenStudy (anonymous):

ohhhh okay!!!

OpenStudy (anonymous):

can you help me with this Explain the additive and multiplicative inverse properties of numbers accurately.

OpenStudy (phi):

in a minute. what is the multiplicative inverse of 1/4 ?

OpenStudy (anonymous):

4/1??

OpenStudy (phi):

yes. but people usually simplify 4 divided by 1. 4/1 = ?

OpenStudy (phi):

1 goes into 4 4 times.

OpenStudy (anonymous):

yeah thats trtue

OpenStudy (phi):

normally you would not bother to write 4/1, you would just write 4 so the multiplicative inverse of 1/4 is 4 and 1/4* 4 = 1 what is the multiplicative inverse of 4 ?

OpenStudy (anonymous):

1/4 right

OpenStudy (phi):

yes. you could either rewrite 4 as 4/1 and flip it to get 1/4 or you could say if 4 is the inverse of 1/4 then 1/4 is the inverse of 4 (they are pairs)

OpenStudy (anonymous):

oh okay!

OpenStudy (phi):

***Explain the additive and multiplicative inverse properties of numbers*** all numbers have an additive inverse. If you are given any number x, then its additive inverse is -1*x a number and its additive inverse *always* add up to zero.

OpenStudy (anonymous):

really i thought additive and milti was hard but is not !

OpenStudy (phi):

If you remember it. If you forget, then it gets hard. all numbers *except zero* have a multiplicative inverse. If you are given any number x (x not equal to zero), then its multiplicative inverse is 1/x the product (i.e. when you multiply them) of a number and its multiplicative inverse is always 1.

OpenStudy (anonymous):

okay

OpenStudy (phi):

the reason you are learning this is that it helps in solving equations in algebra.

OpenStudy (anonymous):

yeah I'min alebra 8th grade keysotne midlle school homeschooled!

OpenStudy (anonymous):

did I do allof this ??? Explain the additive and multiplicative inverse properties of numbers accurately.  Use accurate and appropriate terminology to describe the mathematical ideas presented.  Include at least 5 mathematical examples of the additive inverse, performing any calculations to explain why it is an example of the additive inverse accurately and correctly.  Include at least 5 mathematical examples of the multiplicative inverse, performing any calculations to explain why it is an example of the multiplicative inverse accurately and correctly

OpenStudy (phi):

yes, all of that is answered up above.

OpenStudy (anonymous):

are sure hoe about this Include at least 5 mathematical examples of the additive inverse, performing any calculations to **explain why it is an example of the additive inverse accurately and correctly.***

OpenStudy (anonymous):

Include at least 5 mathematical examples of the multiplicative inverse, performing any calculations to **explain why it is an example of the multiplicative inverse accurately and correctly**

OpenStudy (phi):

if two numbers multiplied together equal 1, then *by definition* they are multiplicative inverses. so if you write \[ \frac{1}{5} \cdot 5 \] and multiply it out to show you get 1 (as you do) that shows that \( \frac{1}{5}\) is the multiplicative inverse of 5 (and also 5 is the multiplicative inverse of \( \frac{1}{5}\) )

OpenStudy (anonymous):

so I have to put that in amy HW

OpenStudy (phi):

You should put in the definition of multiplicative inverse. (and additive inverse) Your notes or book should have the definition (if you copy the definition, be sure to say where you found the definition)

OpenStudy (anonymous):

oh okay soo all my hw is done then

OpenStudy (phi):

after you write it up.

OpenStudy (anonymous):

is this good ?? all numbers have an additive inverse. If you are given any number x, then its additive inverse is -1*x a number and its additive inverse *always* add up to zero. 5+ -5 = 0 7+ -5 = 0 9+ -9 =0 12+ -12=0 14+ -14 = 0 5/8 8/5 6/9 9/6 5/2 2/5 3/6 6/3 7/8 8/7

OpenStudy (phi):

you should try to use your own words (or the book's words, as long as you say, from "title of book") for the bottom part, you should first put in the definition of multiplicative inverse (your words or the books words) then *multiply* each pair and show you get 1 (which proves each pair is a multiplicative inverse)